# Digital currency hardware wallets and the essence of money

### 28 Feb 2026 v4

@@ 3.3 Standards and lobbying
It is plain to see why a hardware manufacturer would lobby in favour of regulations and standards that declare “offline transactions” to be an essential requirement for CBDC. But this argument is unsound: Cash is an excellent offline payment method, and to our knowledge, few if any central banks have suggested that the purpose of CBDC is to replace cash outright. The argument implicitly equivocates between a concept of “offline” that denotes a transaction wherein a user is not required to carry a device with network connectivity, versus a concept of “offline” that denotes a transaction between two counterparties that physically meet in a location without third parties or functioning network infrastructure of any kind nearby. The former scenario is the only one in which users commonly use certified hardware today, for example, single-offline POS payments (e.g. Apple Pay) and network applications that rely upon DRM (e.g. video streaming). Even the specialised Suica card offered by Japan Rail for use at its payment gates, which uses embedded trusted hardware to facilitate payments when connectivity to a central server is broken [Shiibashi 2008], could potentially use locally reachable points of trust to aggregate transaction information rather than assume that network connectivity is either all or nothing. Moreover, the idea that cash could suddenly become unavailable for general use in offline situations (the latter scenario) is an overly simplistic justification for a system-level requirement.
@@
Table 2 summarises the key differences between using a network service as a root of trust and using certified hardware as root of trust. Although certified hardware has the advantage of enabling dual-offline payments, such payments are not actually common in the digital economy. Secure network services can provide the same degree of transaction assurance that secure hardware devices can provide, with the only limitation being that they must be reachable by one of the transacting parties. At the same time, we have argued that because it is feasible for adversarial state actors or organised crime syndicates to compromise individual devices in their possession, it would be too much of a risk for a system relying upon such devices to allow them to spend money that is not linked to attributed accounts, because compromised devices would be able to reuse anonymous money (bearer instruments) without limit. However, with a network service, there is no system-level risk if an end-user device is compromised, since the locus of trust is elsewhere. A network service that provides payment finality222the state in which obligations are discharged between transacting parties. for bearer instruments can be operated by a trusted third party, as in GNU Taler, or it can be oblivious and based on public commitments, so that compromising it does not enable double-spending [Friolo 2025].
@@
Standards that recommend hardware wallets to use a secure element to prevent the exfiltration of private data might be appropriate in the case of transactions involving smartphones, because of their complexity and exposure to malware attacks. However, it is not necessarily the case that all or most digital wallets will be smartphones, despite being the “benchmark in user-accessible features”, as Barresi and Zatti have argued [Barresi 2023]. One reason is price: Although virtually everyone must have a way to engage with the economy, a substantial share of the population cannot afford a smartphone. In 2024, the global median price of a smartphone was US\$92.59, representing 11% of global median per capita income [GDIP 2024]. Another reason is security: with ever more complex operating systems and applications, smartphones are increasingly plagued by security vulnerabilities [NIST], many of which specifically target banking credentials [Fadilpašić 2025]. For a sensitive function like payments, it might be safer and more appropriate to use a dedicated device [Bizama 2025].
@@ 4 Digital currency means tokens, not accounts
Barresi and Zatti assume that wallets would represent end-user money as account balances [Barresi 2023]. However, for ordinary consumers and non-financial businesses to hold digital currency directly, there must be some way for end-users to hold digital currency without the requirement to have an attributed account. Accounts generally involve a relationship between an account holder and a fiduciary, typically a regulated financial institution. A salient property of accounts is that access to assets held within accounts is determined by the identity of the account holder [Auer and Böhme 2020], and the balance of an account represents an entitlement that can be adjudicated, for example, in a court of law. In contrast, tokens are assets that can be held outside of accounts. Tokens can be bearer instruments, meaning that they can be used in a way that does not reference the identity of their holders. Payments can be held offline and transacted online, so the user experience would be similar to an EMV card payment except in which the payer presents digital assets rather than instructions to debit an account.
@@
Some central banks, including the ECB [European Commission 2023], the US Federal Reserve [Federal Reserve 2022], and the Bank of England [Bank of England 2023], have presented arguments that holding limits for wallets are necessary for financial stability and to prevent bank runs. However, as an alternative to the heavy-handed approach of monitoring or restricting wallets, an approach of limiting CBDC withdrawals from banks would be no less effective in preventing bank runs or bank disintermediation than limiting cash withdrawals from banks, and the approach of limiting bank withdrawals from banks has been used successfully in the past, notably in Greece in 2015 [Associated Press 2018]. In this sense, CBDC is no different from cash. To argue that requiring wallets to be managed by parties other than their holders is necessary to address financial stability problems resulting from reduced liquidity is to conflate the technical design requirements for a wallet with economic justifications for eliminating the ability for individuals to hold money outside accounts. However, although some banks and governments might see value in eliminating cash-like instruments and might view CBDC as a vehicle for achieving that objective, such an argument is not commonly articulated.
@@
Importantly, limiting CBDC withdrawals can be done without requiring wallets to implement account-like control mechanisms that undermine the will of their users. Barresi and Zatti suggest that wallets could perform “behavioural analytics” on their users [Barresi 2023]; little could be more disrespectful of a user’s autonomy. Similarly, the argument that digital currency could be used as a way to implement “programmable money” that can only be spent in pre-defined ways is a non-starter. Such a mechanism would not only undermine the singleness of money but also open a Pandora’s box of ethical concerns [Li 2025], some of which have already been realised by the failed cashless debit card scheme in Australia [Greenacre 2023]. However, banks could impose simple rules to limit the aggregate size of withdrawals that can be made from their accounts into CBDC, both in a single transaction and over a period of time, and such rules could be potentially combined with identity information to limit the amount that an end-user could withdraw throughout the system. Such rules could also be combined with periodic token expiry, as a mechanism to prevent hoarding [Goodell 2021]. Periodic token expiry is less intrusive than mechanisms that restrict or control wallets directly, and it can be made transparent to users by implementing a mechanism for wallets to automatically refresh any assets that are close to the expiration time.
@@
Although a third party is necessary for fair exchange (and indeed trusted hardware or closed-source software blobs can act as this third party de facto), Pagnia and Gärtner have demonstrated that the role of the third party can be performed “in zero-knowledge”. Specifically, a third party can facilitate final settlement of an asset transfer involving two transacting parties, without having a custodial relationship with either party, without knowing the identity of either party, without knowing how much is being transferred, and even without knowing that tokens are being transferred at all [Pagnia 1999]. Importantly, with oblivious management, an issuer does not rely upon a central ledger to track tokens individually, although the issuer must trust that the system that collects transactions and produces rolled-up aggregation entries, one per time period, will not equivocate. This trust can be enhanced by requiring that system to externalise its commitment to a single history, for example by sending some or all of its aggregation entries to the issuer, or by writing them to a public bulletin board or distributed ledger.
@@
Digital assets can indeed be held by users without accounts of any kind, just like cash and other physical assets. An example of a digital asset that can be held without an account is an e-cash token in Chaum’s model [Chaum 1982, Chaum 2021], which requires every transaction to function as a redemption, and it has also been shown that digital assets can be transferred without redemption [Goodell 2023, Friolo 2025]. Accounts for recipients might be desirable, but they are not technically necessary. As described above, it has been proven that the prevention of double-spending does not imply the existence of a traceable record of all the holdings and movements through sequential time. In Chaum’s original proposal, an asset can be transferred to a recipient without revealing information about the payer, and, unlike oblivious management, double-spending is prevented by maintaining a record of the spent tokens. In the general case, and in the approach taken by oblivious management, to prevent double-spending, it is sufficient to register non-repudiable commitments from payers, which may take the form of a hash or zero-knowledge proof that can be verified without knowledge of the transaction [Pagnia 1999], and proofs of these commitments can be delivered along with the asset data rather than centrally recorded [Goodell 2023, Friolo 2025]. Oblivious management presupposes that issuers can be held responsible for what they have issued and that proofs of commitments that cryptographically demonstrate the validity and integrity of an asset constitute sufficient evidence of an issuer’s responsibility to honour a redemption. As discussed earlier, specific payment architectures that implement these mechanisms have been proposed.
Tokens are also important for privacy. Systems that make use of transactions involving balances, even if the transactions themselves are privacy-preserving [Wüst 2022, Kiayias 2022], intrinsically require end-users to retain data produced during a transaction that can be used to prove that the transaction was carried out according to the rules. This means that transactions are not independent of one another. The property of transaction independence is important, because retaining data related to past transactions introduces the risk that the user could be de-anonymised via poor operational security or blackmailed to reveal information about those transactions [Friolo 2025]. Because it must be possible for balances to be tracked through all of the movements of that balance, transactions in systems that represent money as balances cannot achieve transaction independence. However, systems that implement money as tokens are not subject to this constraint.
@@ 5 Separating issuance from transaction processing
Even if transactions are done with network access to a mutually trusted third party, there is no reason to assume (a) that the third party must be the issuer, or (b) that the issuer must be involved in transaction processing in real-time. Chaum’s original proposal assumed both conditions [Chaum 1982], but they are not requirements in general. In particular, with oblivious management, it is possible to register a (zero-knowledge) commitment to transfer a digital asset without communicating directly with the issuer. For a transaction to have finality, assuming every transfer specifies the payee with a unique single-use key, it is sufficient for both parties have confidence in a mutually trusted third party, acting as a notary, to eventually post the commitment (or an aggregation that comprehends the commitment) to an append-only bulletin board that maps single-use keys to commitments and that is trusted by the issuer [Goodell 2023]. The append-only bulletin board can be operated by the issuer directly, by a party that the issuer trusts, or as a distributed ledger operated by a set of parties that the issuer trusts in aggregate. Provided that the operator of the bulletin board does not equivocate, then since no two tokens can be transferred to the same single-use key and have a valid proof of provenance, it is not necessary for the issuer to maintain a ledger that records the registration status of individual tokens [Goodell 2023].
@@
Even if we assume that all tokens are issued by a central issuer, such as a central bank, it is certainly possible to use a decentralised system to conduct transactions using those tokens, and wallets can be designed to respect this possibility. Trusting a central bank (or other issuing authority, such as one that issues e-money) to issue currency and maintain its integrity is certainly different from trusting it to process e-cash transactions. In particular, it is possible to use a decentralised system to transact centrally issued money. Just as cash can be managed with a two-tiered approach, e-cash can also be managed with a two-tiered approach. If we assume, following Chaum, that individual e-cash assets are generated for one-time use by consumers, then a central bank can provide private-sector banks with “genesis” assets that can be burned, pari passu, by private-sector banks in the process of creating single-use e-cash tokens for consumers. Then, the single-use assets can be recirculated within the system by private-sector banks indefinitely via the same process before ultimately being redeemed with the central bank [Friolo 2025]. The use of a secure element to assure trust seems to implicitly require the transacting parties to respect the authority of a specific hardware manufacturer, and the power implicit to this authority might be incompatible with the concept of decentralisation if all transacting parties are forced to use secure elements produced by a limited set of manufacturers. To require the use of a hardware root of trust with remote attestation unnecessarily limits decentralisation.
@@ 6 Wallets as containers for bearer instruments
The case for trusted hardware, as for other core mechanisms of proposed payment systems, involves the requirement to ensure the integrity of money when counterparties lack access to a mutually trusted third party. In principle, trusted hardware can act as a trusted third party to enforce system-level rules on user devices directly, by implementing rules that prevent digital assets held in a wallet from being spent more than once. However, considering the security risks associated with compromised devices, it is better to enforce such rules over the network instead, if it is possible to do so. This can be done, for example, by requiring the issuer to validate all transactions, as proposed by Chaum and others [Chaum 1982, Chaum 2021]. However, there is reason to believe the costs, risks, and public accountability concerns associated with having a single central operator perform issuance, settlement, and compliance functions might be too great for the CBDC use case [Goodell 2021].
@@
The digital currency design proposed by Goodell et al [Goodell 2023] and further elaborated by Friolo et al [Friolo 2025] offers an alternative approach to achieving scalability, replacing global consensus about the status and ownership of individual assets, and the related costs of such consensus common to modern cryptocurrency systems, with a combination of local trust and a mechanism for keeping local trust providers honest. By integrating a fungible token created using Chaum’s method that delivers privacy by design for consumers into a non-fungible asset that can be transferred with strong traceability following the initial creation, it enables the separation of the costs and risks associated with the aforementioned roles that enable payments. With this design, the third parties that provide transaction finality do not manage tokens, the issuer is not involved in individual transactions, and regulated entities, such as banks, can continue to provide compliance enforcement (e.g. income monitoring for the purposes of taxation and sanctions enforcement) as a condition of exchanging a spent token for a fresh token or another form of money.
@@
A much better way to address the limited set of cases in which “fully offline” payments (that is, those where neither the payer nor the payee have access to any suitable third parties or network infrastructure) are necessary would be to support the use of cash itself where cash continues to work well, such as in remote locations and environments with low connectivity, while avoiding all of the aforementioned problems associated with trusted hardware. To address the observed reduction of cash acceptance in urban environments [Mansuri and Vyas 2024, De Nederlandsche Bank 2025] by providing digital currency, whilst encouraging the provision of facilities for converting between cash and digital currency at the boundary of “unbuilt” environments, would be an entirely legitimate approach. In particular, if one potential benefit of CBDC in an age of waning cash infrastructure [^4] Evidence for the decline of cash is growing [Access to Cash Review 2019, Tisher 2020], and it is acknowledged as an important risk by central banks, including the ECB [European Commission 2023a]. is the ability to electronically transport money to remote population centres during crises [Goodell 2020], then there is no reason to require that central bank money distributed in this manner must remain digital. For example, service providers in a particular locality can provide physical tokens in exchange for CBDC, where the CBDC is transferred using one of the methods described earlier, much as a casino can provide chips in exchange for cash or bank money.

### 12 Mar 2026 v5

@@ 3.3 Standards and lobbying
It is plain to see why a hardware manufacturer would lobby in favour of regulations and standards that declare “offline transactions” to be an essential requirement for CBDC. But this argument is unsound: Cash is an excellent offline payment method, and to our knowledge, few if any central banks have suggested that the purpose of CBDC is to replace cash outright. The argument implicitly equivocates between a concept of “offline” that denotes a transaction wherein a user is not required to carry a device with network connectivity, versus a concept of “offline” that denotes a transaction between two counterparties that physically meet in a location without third parties or functioning network infrastructure of any kind nearby. The former scenario is the only one in which users commonly use certified hardware today, for example, single-offline POS payments (e.g. Apple Pay) and network applications that rely upon DRM (e.g. video streaming). Even the use case for the specialised Suica card, which is offered by Japan Rail for use at its payment gates and uses embedded trusted hardware to facilitate payments when connectivity to a central server is broken [Shiibashi 2008], could potentially leverage locally reachable points of trust to aggregate transaction information rather than assume that network connectivity is either all or nothing. Moreover, the idea that cash could suddenly become unavailable for general use in offline situations (the latter scenario) is an overly simplistic justification for a system-level requirement.
@@
Table 2 summarises the key differences between using a network service as a root of trust and using certified hardware as root of trust. Although certified hardware has the advantage of enabling dual-offline payments, such payments are not actually common in the digital economy. Secure network services can provide the same degree of transaction assurance that secure hardware devices can provide, with the only limitation being that they must be reachable by one of the transacting parties. At the same time, we have argued that because it is feasible for adversarial state actors or organised crime syndicates to compromise individual devices in their possession, it would be too much of a risk for a system relying upon such devices to allow them to spend money that is not linked to attributed accounts, because compromised devices would be able to reuse anonymous money (bearer instruments) without limit. However, with a network service, there is no system-level risk if an end-user device is compromised, since the locus of trust is elsewhere. A network service that provides payment finality222the state in which obligations are discharged between transacting parties. for bearer instruments can be operated by a trusted third party, as in GNU Taler, or it can be oblivious and based on public commitments, so that compromising it does not enable double-spending [Friolo et al. 2025].
@@
Standards that recommend hardware wallets to use a secure element to prevent the exfiltration of private data might be appropriate in the case of transactions involving smartphones, because of their complexity and exposure to malware attacks. However, it is not necessarily the case that all or most digital wallets will be smartphones, despite being the “benchmark in user-accessible features”, as Barresi and Zatti have argued [Barresi and Zatti 2023]. One reason is price: Although virtually everyone must have a way to engage with the economy, a substantial share of the population cannot afford a smartphone. In 2024, the global median price of a smartphone was US\$92.59, representing 11% of global median per capita income [GDIP 2024]. Another reason is security: with ever more complex operating systems and applications, smartphones are increasingly plagued by security vulnerabilities [NIST], many of which specifically target banking credentials [Fadilpašić 2025]. For a sensitive function like payments, it might be safer and more appropriate to use a dedicated device [Bizama et al. 2025].
@@ 4 Digital currency means tokens, not accounts
Barresi and Zatti assume that wallets would represent end-user money as account balances [Barresi and Zatti 2023]. However, for ordinary consumers and non-financial businesses to hold digital currency directly, there must be some way for end-users to hold digital currency without the requirement to have an attributed account. Accounts generally involve a relationship between an account holder and a fiduciary, typically a regulated financial institution. A salient property of accounts is that access to assets held within accounts is determined by the identity of the account holder [Auer and Böhme 2020], and the balance of an account represents an entitlement that can be adjudicated, for example, in a court of law. In contrast, tokens are assets that can be held outside of accounts. Tokens can be bearer instruments, meaning that they can be used in a way that does not reference the identity of their holders. Payments can be held offline and transacted online, so the user experience would be similar to an EMV card payment except in which the payer presents digital assets rather than instructions to debit an account.
@@
Some central banks, including the ECB [European Commission 2023], the US Federal Reserve [Federal Reserve 2022], and the Bank of England [Bank of England and HM Treasury 2023], have presented arguments that holding limits for wallets are necessary for financial stability and to prevent bank runs. However, as an alternative to the heavy-handed approach of monitoring or restricting wallets, an approach of limiting CBDC withdrawals from banks would be no less effective in preventing bank runs or bank disintermediation than limiting cash withdrawals from banks, and the approach of limiting bank withdrawals from banks has been used successfully in the past, notably in Greece in 2015 [Associated Press 2018]. In this sense, CBDC is no different from cash. To argue that requiring wallets to be managed by parties other than their holders is necessary to address financial stability problems resulting from reduced liquidity is to conflate the technical design requirements for a wallet with economic justifications for eliminating the ability for individuals to hold money outside accounts. However, although some banks and governments might see value in eliminating cash-like instruments and might view CBDC as a vehicle for achieving that objective, such an argument is not commonly articulated.
@@
Importantly, limiting CBDC withdrawals can be done without requiring wallets to implement account-like control mechanisms that undermine the will of their users. Barresi and Zatti suggest that wallets could perform “behavioural analytics” on their users [Barresi and Zatti 2023]; little could be more disrespectful of a user’s autonomy. Similarly, the argument that digital currency could be used as a way to implement “programmable money” that can only be spent in pre-defined ways is a non-starter. Such a mechanism would not only undermine the singleness of money but also open a Pandora’s box of ethical concerns [Li and Shen 2025], some of which have already been realised by the failed cashless debit card scheme in Australia [Greenacre 2023]. However, banks could impose simple rules to limit the aggregate size of withdrawals that can be made from their accounts into CBDC, both in a single transaction and over a period of time, and such rules could be potentially combined with identity information to limit the amount that an end-user could withdraw throughout the system. Such rules could also be combined with periodic token expiry, as a mechanism to prevent hoarding [Goodell et al. 2021]. Periodic token expiry is less intrusive than mechanisms that restrict or control wallets directly, and it can be made transparent to users by implementing a mechanism for wallets to automatically refresh any assets that are close to the expiration time.
@@
Although a third party is necessary for fair exchange (and indeed trusted hardware or closed-source software blobs can act as this third party de facto), Pagnia and Gärtner have demonstrated that the role of the third party can be performed “in zero-knowledge”. Specifically, a third party can facilitate final settlement of an asset transfer involving two transacting parties, without having a custodial relationship with either party, without knowing the identity of either party, without knowing how much is being transferred, and even without knowing that tokens are being transferred at all [Pagnia and Gärtner 1999]. Importantly, with oblivious management, an issuer does not rely upon a central ledger to track tokens individually, although the issuer must trust that the system that collects transactions and produces rolled-up aggregation entries, one per time period, will not equivocate. This trust can be enhanced by requiring that system to externalise its commitment to a single history, for example by sending some or all of its aggregation entries to the issuer, or by writing them to a public bulletin board or distributed ledger.
@@
Digital assets can indeed be held by users without accounts of any kind, just like cash and other physical assets. An example of a digital asset that can be held without an account is an e-cash token in Chaum’s model [Chaum 1983, Chaum et al. 2021], which requires every transaction to function as a redemption, and it has also been shown that digital assets can be transferred without redemption [Goodell et al. 2023, Friolo et al. 2025]. Accounts for recipients might be desirable, but they are not technically necessary. As described above, it has been proven that the prevention of double-spending does not imply the existence of a traceable record of all the holdings and movements through sequential time. In Chaum’s original proposal, an asset can be transferred to a recipient without revealing information about the payer, and, unlike oblivious management, double-spending is prevented by maintaining a record of the spent tokens. In the general case, and in the approach taken by oblivious management, to prevent double-spending, it is sufficient to register non-repudiable commitments from payers, which may take the form of a hash or zero-knowledge proof that can be verified without knowledge of the transaction [Pagnia and Gärtner 1999], and proofs of these commitments can be delivered along with the asset data rather than centrally recorded [Goodell et al. 2023, Friolo et al. 2025]. Oblivious management presupposes that issuers can be held responsible for what they have issued and that proofs of commitments that cryptographically demonstrate the validity and integrity of an asset constitute sufficient evidence of an issuer’s responsibility to honour a redemption. As discussed earlier, specific payment architectures that implement these mechanisms have been proposed.
Tokens are also important for privacy. Systems that make use of transactions involving balances, even if the transactions themselves are privacy-preserving [Wüst et al. 2022, Kiayias et al. 2022], intrinsically require end-users to retain data produced during a transaction that can be used to prove that the transaction was carried out according to the rules. This means that transactions are not independent of one another. The property of transaction independence is important, because retaining data related to past transactions introduces the risk that the user could be de-anonymised via poor operational security or blackmailed to reveal information about those transactions [Friolo et al. 2025]. Because it must be possible for balances to be tracked through all of the movements of that balance, transactions in systems that represent money as balances cannot achieve transaction independence. However, systems that implement money as tokens are not subject to this constraint.
@@ 5 Separating issuance from transaction processing
Even if transactions are done with network access to a mutually trusted third party, there is no reason to assume (a) that the third party must be the issuer, or (b) that the issuer must be involved in transaction processing in real-time. Chaum’s original proposal assumed both conditions [Chaum 1983], but they are not requirements in general. In particular, with oblivious management, it is possible to register a (zero-knowledge) commitment to transfer a digital asset without communicating directly with the issuer. For a transaction to have finality, assuming every transfer specifies the payee with a unique single-use key, it is sufficient for both parties have confidence in a mutually trusted third party, acting as a notary, to eventually post the commitment (or an aggregation that comprehends the commitment) to an append-only bulletin board that maps single-use keys to commitments and that is trusted by the issuer [Goodell et al. 2023]. The append-only bulletin board can be operated by the issuer directly, by a party that the issuer trusts, or as a distributed ledger operated by a set of parties that the issuer trusts in aggregate. Provided that the operator of the bulletin board does not equivocate, then since no two tokens can be transferred to the same single-use key and have a valid proof of provenance, it is not necessary for the issuer to maintain a ledger that records the registration status of individual tokens [Goodell et al. 2023].
@@
Even if we assume that all tokens are issued by a central issuer, such as a central bank, it is certainly possible to use a decentralised system to conduct transactions using those tokens, and wallets can be designed to respect this possibility. Trusting a central bank (or other issuing authority, such as one that issues e-money) to issue currency and maintain its integrity is certainly different from trusting it to process e-cash transactions. In particular, it is possible to use a decentralised system to transact centrally issued money. Just as cash can be managed with a two-tiered approach, e-cash can also be managed with a two-tiered approach. If we assume, following Chaum, that individual e-cash assets are generated for one-time use by consumers, then a central bank can provide private-sector banks with “genesis” assets that can be burned, pari passu, by private-sector banks in the process of creating single-use e-cash tokens for consumers. Then, the single-use assets can be recirculated within the system by private-sector banks indefinitely via the same process before ultimately being redeemed with the central bank [Friolo et al. 2025]. The use of a secure element to assure trust seems to implicitly require the transacting parties to respect the authority of a specific hardware manufacturer, and the power implicit to this authority might be incompatible with the concept of decentralisation if all transacting parties are forced to use secure elements produced by a limited set of manufacturers. To require the use of a hardware root of trust with remote attestation unnecessarily limits decentralisation.
@@ 6 Wallets as containers for bearer instruments
The case for trusted hardware, as for other core mechanisms of proposed payment systems, involves the requirement to ensure the integrity of money when counterparties lack access to a mutually trusted third party. In principle, trusted hardware can act as a trusted third party to enforce system-level rules on user devices directly, by implementing rules that prevent digital assets held in a wallet from being spent more than once. However, considering the security risks associated with compromised devices, it is better to enforce such rules over the network instead, if it is possible to do so. This can be done, for example, by requiring the issuer to validate all transactions, as proposed by Chaum and others [Chaum 1983, Chaum et al. 2021]. However, there is reason to believe the costs, risks, and public accountability concerns associated with having a single central operator perform issuance, settlement, and compliance functions might be too great for the CBDC use case [Goodell et al. 2021].
@@
The digital currency design proposed by Goodell et al [Goodell et al. 2023] and further elaborated by Friolo et al [Friolo et al. 2025] offers an alternative approach to achieving scalability, replacing global consensus about the status and ownership of individual assets, and the related costs of such consensus common to modern cryptocurrency systems, with a combination of local trust and a mechanism for keeping local trust providers honest. By integrating a fungible token created using Chaum’s method that delivers privacy by design for consumers into a non-fungible asset that can be transferred with strong traceability following the initial creation, it enables the separation of the costs and risks associated with the aforementioned roles that enable payments. With this design, the third parties that provide transaction finality do not manage tokens, the issuer is not involved in individual transactions, and regulated entities, such as banks, can continue to provide compliance enforcement (e.g. income monitoring for the purposes of taxation and sanctions enforcement) as a condition of exchanging a spent token for a fresh token or another form of money.
@@
A much better way to address the limited set of cases in which “fully offline” payments (that is, those where neither the payer nor the payee have access to any suitable third parties or network infrastructure) are necessary would be to support the use of cash itself where cash continues to work well, such as in remote locations and environments with low connectivity, while avoiding all of the aforementioned problems associated with trusted hardware. To address the observed reduction of cash acceptance in urban environments [Mansuri and Vyas 2024, De Nederlandsche Bank 2025] by providing digital currency, whilst encouraging the provision of facilities for converting between cash and digital currency at the boundary of “unbuilt” environments, would be an entirely legitimate approach. In particular, if one potential benefit of CBDC in an age of waning cash infrastructure [^4] Evidence for the decline of cash is growing [Access to Cash Review 2019, Tisher et al. 2020], and it is acknowledged as an important risk by central banks, including the ECB [European Commission 2023a]. is the ability to electronically transport money to remote population centres during crises [Goodell et al. 2020], then there is no reason to require that central bank money distributed in this manner must remain digital. For example, service providers in a particular locality can provide physical tokens in exchange for CBDC, where the CBDC is transferred using one of the methods described earlier, much as a casino can provide chips in exchange for cash or bank money.